OpenStudy (anonymous):

Help please.Evaluate. lim x-0 {(sin2x +3x)/( 2x +sin3x)] Please do it with steps

OpenStudy (australopithecus):

You can use identities

OpenStudy (australopithecus):

is that sin(2x) or sin^(2)(x)

OpenStudy (anonymous):

sin2x

OpenStudy (anonymous):

Im confused, please make me see your steps

OpenStudy (australopithecus):

yeah I guess you could use L'H rule

OpenStudy (anonymous):

Let me reformulate:

OpenStudy (australopithecus):

Take the derivative of the top and the bottom. Treat the neumorator and denominator as two seperate non fractional equations

OpenStudy (australopithecus):

to use L'H rule

OpenStudy (australopithecus):

lim x-0 {(sin2x +3x)/( 2x +sin3x)] = L'H = lim (2cos(x) + 3)/(2 + 3cos(x)) x->0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}[(\sin{2x} + 3x)/(2x + \sin{3x})]\]By L.H., we get:\[\lim_{x \rightarrow 0}[(2\cos{2x} + 3)/(3\cos{3x} + 2)]\]Try it out now

OpenStudy (australopithecus):

= 5/5 = 1

OpenStudy (australopithecus):

If he isn't allowed to use L'H rule perhaps he is better of using trig Identities

OpenStudy (anonymous):

Indeed, but they are a pain to use on an exam, except for the identity one with sin^2 + cos^2, :-)

OpenStudy (australopithecus):

sorry I mean lim ln(x)ln(1/x) x->0+

OpenStudy (anonymous):

I have done this ..PLEASE PLEASE PLEASE VERIFY MY STEPS, TELL ME IF IM RIGHT OR WRONG

OpenStudy (australopithecus):

If you want to use trig identites rememebr that sin(2x) = 2sin(x)cos(x) I can't recall what the sin(x/y) -

OpenStudy (australopithecus):

no right sorry :)

OpenStudy (australopithecus):

7/2 / 7/2 = 1

OpenStudy (anonymous):

M i WRONG?

OpenStudy (australopithecus):

No you got 1 so you are right I have no idea what you did but you got the right answer

OpenStudy (anonymous):

thats wht I was telling PLEASE CHECK MY STEPS

OpenStudy (australopithecus):

You need to use Identites sin(2x) = 2sin(x)cos(x) then there is another identity that is something like sin(xy) = sin(x)cos(y) - sin(x)cos(y)

OpenStudy (australopithecus):

These are the identites you need to use sin(A+B) = sinAcosB + cosAsinB sin(2x) = 2sin(x)cos(x) To solve this sin(3x) = sin(2x + x)

OpenStudy (anonymous):

or we can do like this lim x-0 {(sin2x)/x +3}/ {2 +(sin3x)/x} put sin2x/x = (2sin2x)/2x = 2and (3sin3x)/3x =3bcz lim x-0 (sinx)/x so we will get (2+3)/3+2) = 1

OpenStudy (anonymous):

@rahul33333 Almost right, but d(sin3x)/dx = 3sin(3x) not sin(3x)/3. Other than that, it's correct what you did.

OpenStudy (anonymous):

I mean the same for cos(2x)

OpenStudy (anonymous):

ohhhhooo, I did a mistake yes yes. thanks , MEANS I LEARN L HOSPITAL RULE !!! wow

OpenStudy (anonymous):

Kudos :-)

OpenStudy (australopithecus):

Crud you need also this identity cos(2x) = cos²(x) - sin²(x) we can get rid of those ugly trig functions by (2cos(x)sin(x) +3x)/(2x + sin(2x)cos(x) +sin(x)cos(2x))) = (2cos(x)sin(x) + 3x)/(2x + 2sin(x)cos(x)cos(x) + sin(x)(cos^(2)(x) - sin^(2)(x)))

OpenStudy (anonymous):

fromer was a better way, anyway thanks

OpenStudy (australopithecus):

lol have fun simplifying that though :)